# Flexibility Pricing in Distribution Systems: A Direct Method Aligned with Flexibility Models

### v1

@@ II-B2 Battery Energy Storage Systems
The flexibility of a BESS lies in its ability to charge and discharge within the battery’s capacity. The energy level of a BESS typically needs periodic balancing to preserve sufficient capacity for potential charging and discharging demands. Therefore, the inconvenience caused to a BESS user by flexibility activation can be quantified by the unbalanced energy at times of balancing.
@@ II-B3 Heat Pumps
We define the baseline power and accumulated energy consumption profiles to always maintain the indoor temperature at the set point $\theta^{\text{set}}_{t},\forall t\in[T]$. Activating the flexibility of HPs leads to the fluctuation of indoor temperature within the upward and downward adjustment ranges relative to the temperature set point, denoted by $\Delta\hat{\theta}_{t}$ and $\Delta\check{\theta}_{t}$, respectively, (with vector forms $\Delta\hat{\boldsymbol{\theta}}\in\mathbb{R}^{T}_{+}$ and $\Delta\check{\boldsymbol{\theta}}\in\mathbb{R}^{T}_{+}$). Regarding the relationship between the temperature adjustment ranges, $\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\boldsymbol{\theta}}$, and the upward and downward adjustment ranges in power, $\Delta\hat{\mathbf{p}}$ and $\Delta\check{\mathbf{p}}$, and accumulated energy consumption, $\Delta\hat{\mathbf{e}}$ and $\Delta\check{\mathbf{e}}$, the following Proposition holds:
@@
The upward and downward temperature adjustment ranges, $\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\boldsymbol{\theta}}$, can be linearly transformed into the upward and downward adjustment ranges in accumulated energy consumption, $\Delta\hat{\mathbf{e}}$ and $\Delta\check{\mathbf{e}}$, respectively, by:
$\Delta\hat{\mathbf{e}}={\mathbf{D}}\Delta\hat{\boldsymbol{\theta}},\;\;\Delta\check{\mathbf{e}}={\mathbf{D}}\Delta\check{\boldsymbol{\theta}},$ (6)
where ${\mathbf{D}}$ is a $T\times T$ invertible matrix, defined as:
${\mathbf{D}}\triangleq\Delta_{T}\cdot\frac{H}{\eta}\left[{\begin{array}[]{cccc}1&{}\hfil&{}\hfil&{}\hfil\\ {1-\alpha}&1&{}\hfil&{}\hfil\\ \vdots&\ddots&\ddots&{}\hfil\\ {1-\alpha}&\cdots&{1-\alpha}&1\end{array}}\right].$
However, $\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\boldsymbol{\theta}}$ cannot be respectively transformed into upward and downward power adjustment ranges $\Delta\hat{\mathbf{p}}$ and $\Delta\check{\mathbf{p}}$.
Proof: See Appendix A.
The rationale behind Proposition 1 can be intuitively understood as follows: the indoor temperature is primarily influenced by the accumulated power over time rather than the power in each time slot itself. The transformation in (6) allows aggregators to have a contract with HP users specifying the compensation for per unit temperature adjustment range and to transform this compensation into cost coefficients of energy adjustment range. We denote by $\hat{\boldsymbol{\rho}}\in\mathbb{R}^{T}_{+}$ and $\check{\boldsymbol{\rho}}\in\mathbb{R}^{T}_{+}$ the vectors of compensations for the activated upward and downward temperature adjustment ranges, respectively. Then, the total cost of the activated flexibility of an HP is:
$C_{\text{HP}}=\hat{\boldsymbol{\rho}}^{\top}\Delta\hat{\boldsymbol{\theta}}+\check{\boldsymbol{\rho}}^{\top}\Delta\check{\boldsymbol{\theta}},$
@@
$C_{\text{HP}}=\hat{\boldsymbol{\rho}}^{\top}\mathbf{D}^{-1}\Delta\hat{\mathbf{e}}+\check{\boldsymbol{\rho}}^{\top}\mathbf{D}^{-1}\Delta\check{\mathbf{e}},$
@@
Fig. 4 illustrates adjustment ranges in the indoor temperature, the corresponding adjustment ranges in accumulated energy consumption, and cost coefficients. Cost coefficients of energy adjustment ranges shown in this figure, which increase with time, are generated based on a constant cost coefficient of temperature adjustment range across the time horizon. Take $\check{\mathbf{c}}_{\text{e}}$ as an example, assuming that $\check{\rho}_{t}=\check{\rho}_{0},\forall t\in[T]$, then $\check{\mathbf{c}}_{\text{e}}=([\check{\rho}_{0},\check{\rho}_{0},\cdots,\check{\rho}_{0}]\mathbf{D}^{-1})^{\top}=\frac{\eta}{H\Delta_{T}}\check{\rho}_{0}[{\alpha^{T-1}},...,{\alpha^{2}},\alpha,1]^{\top}$. As $10<\alpha<$ 1 by definition, $\check{c}_{\text{e},t}$ increases with the time index $t$. Besides, $\check{\mathbf{c}}_{\text{e}}$ is typically larger than $\hat{\mathbf{c}}_{\text{e}}$ because downward temperature adjustments cause more discomfort to the people inside the building than upward adjustments considering the primary task of the HP is heating.
@@
Finally, we note that $\hat{\boldsymbol{\rho}}$ and $\check{\boldsymbol{\rho}}$ should not be uniform across all buildings, as the thermal dynamic parameters and HPs’ COPs vary. Identical temperature adjustments can lead to different energy adjustments in different buildings, so applying uniform $\hat{\boldsymbol{\rho}}$ and $\check{\boldsymbol{\rho}}$ would be unfair to HP users in a well-insulated building. Thus, $\hat{\boldsymbol{\rho}}$ and $\check{\boldsymbol{\rho}}$ should be customized based on the thermal dynamic parameters of the building and the HP’s COP. We suggest setting $\hat{\boldsymbol{\rho}}$ and $\check{\boldsymbol{\rho}}$ proportional to $\frac{H\Delta_{T}}{\eta}$, which aligns the HP’s flexibility cost coefficients with the contribution to energy adjustment ranges.
III Model and Cost for Aggregated Flexibility
We have discussed the flexibility cost formulation for individual DERs. This section focuses on the aggregator’s perspective, namely the model and cost for the aggregated flexibility, which provides the bidding information of aggregators in the DSO flexibility market. We denote the set of DERs in the aggregator as $\mathcal{K}$, indexed by $k$, and the power vector of the $k$-th DER as $\mathbf{p}_{k}\in\mathbb{R}^{T}$. Hence, the aggregated power is $\mathbf{P}\triangleq\sum\nolimits_{k\in\mathcal{K}}\mathbf{p}_{k}$. The aggregated flexibility model refers to the feasible region of $\mathbf{P}$. Given the individual flexibility constraints (2)-(4), the aggregated flexibility, as derived in our previous works [15] and [21], can be formulated as follows:
@@ III Model and Cost for Aggregated Flexibility
where $\mathbf{U}$ is an ${M\times T}$ coefficient matrix with each row vector being a ${1\times T}$ vector composed of 0s and 1s, and $M$ represents the number of constraints that is dependent on whether an exact or an approximate model is used. Constraints (7) define a polytope in the $\mathbb{R}^{T}$ space with a specific shape. We let $\mathbb{B}\triangleq\{0,1\}$ and $\mathbb{B}^{T}$ denote the the set of all $T$-dimensional row vectors composed solely of 0s and 1s. If the row vectors of $\mathbf{U}$ span $\mathbb{B}^{T}$ except for $\mathbf{0}$, i.e., $M=2^{T}-1$, $\mathbf{U}$ corresponds to the exact aggregation model. When the row vector set of $\mathbf{U}$ is a subset of $\mathbb{B}^{T}$, $\mathbf{U}$ corresponds to an approximate model. Bound parameters $\underline{\boldsymbol{\phi}}\in\mathbb{R}^{M}$ and $\overline{\boldsymbol{\phi}}\in\mathbb{R}^{M}$ are derived from all the individual power and energy bound parameters in (2)-(3). Reducing the number of rows of $\mathbf{U}$ creates an approximate model with low complexity, and appropriately adjusting $\underline{\boldsymbol{\phi}}$ and $\overline{\boldsymbol{\phi}}$ forms either an outer or inner approximation. For example, setting the coefficient matrix $\mathbf{U}$ to only include power and accumulated energy constraints and the bound parameters $\underline{\boldsymbol{\phi}}$ and $\overline{\boldsymbol{\phi}}$ to the sum of all individual power and energy bounds yields an outer approximation [15], and properly shrinking the gap between $\underline{\boldsymbol{\phi}}$ and $\overline{\boldsymbol{\phi}}$ forms an inner approximation [21].
We are now ready to introduce the flexibility cost formulation for the aggregated flexibility model (7). For the sake of clarity, we hereafter refer to the product of $\mathbf{U}$ and the aggregated power profile $\mathbf{P}$, i.e., $\mathbf{UP}$, as the flexibility trajectory and the product of $\mathbf{U}$ and the aggregated baseline power profile $\mathbf{UP}^{\text{base}}$, i.e., $\mathbf{UP}^{\text{base}}$, as the baseline flexibility trajectory. Similar to the flexibility model (2)-(4) for individual DERs, activating the aggregated flexibility means that the flexibility trajectory $\mathbf{UP}$ can deviate within a certain range relative to the baseline trajectory $\mathbf{UP}^{\text{base}}$, which can be formulated as:
@@
where $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ in +$\mathbb{R}^{M}_{+}$ denote the upward and downward cost coefficients for flexibility adjustment ranges, respectively, which the aggregator bids to the DSO market.
The next question is how the aggregator can calculate the cost coefficients $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$. Principally, $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ are related to the aggregator’s cost of purchasing flexibility from the DER users and its operation costs. Aggregators may develop various strategies to determine $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ for bidding in the DSO flexibility market. Since the focus of this paper is primarily on the formulation of flexibility costs, we here use a simple approach to estimate $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ without a detailed discussion of the bidding strategies of aggregators.
First, it should be noted that the flexibility model (7) includes constraints of power and accumulated energy consumption of the aggregator. In other words, the power and accumulated energy consumption are a part of the flexibility trajectory $\mathbf{UP}$. The power and accumulated energy part of cost coefficients $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ are estimated via a weighted average method, with the weights being the ratio of the maximum power and energy adjustment ranges of each DER, i.e.,
@@
If the aggregated flexibility model (7) includes constraints beyond power and accumulated energy consumption, i.e., $M>|\mathcal{I}_{\text{p,e}}|$, the cost coefficients of adjustment ranges in those additional parts should be directly set to zero. This setting avoids the duplication of the cost calculation, as power and accumulated energy adjustments could lead to adjustments in other segments of the flexibility trajectory. This does not imply that other segments of the flexibility trajectory do not contribute to the flexibility cost of the aggregator since those segments are subject to constraints and indirectly influence the aggregator’s total flexibility cost $C_{\text{agg}}$. Indeed, for simplicity, the DSO can require aggregators to model their flexibility using only power and accumulated energy constraints and to submit the cost coefficients only for adjustment ranges in this segment. Nevertheless, since we have designed a more general flexibility cost formulation, it holds the potential to be applied to DSO markets with more accurate flexibility models.
IV DSO Flexibility Market Clearing and Marginal Flexibility Prices
We consider the scenario where the DSO procures flexibility from aggregators for energy arbitrage and reserve service provision at the transmission level. The DSO flexibility market should 1) determine the reference power profile and reserve capacities offered to the transmission system, 2) determine the activated adjustment range for the flexibility trajectory of each aggregator within the distribution system, and 3) distribute the revenues obtained from the transmission level to each aggregator while ensuring incentive compatibility, revenue adequacy, and a non-profit DSO.
IV-A Modeling of the DSO Flexibility Market Clearing
The objective of the DSO flexibility market clearing is to maximize social welfare, which is equivalent to minimizing the net cost of the distribution system in this paper, formulated as:

### v4

@@ II-B2 Battery Energy Storage Systems
The flexibility of a BESS lies in its ability to charge and discharge within the battery’s capacity. The energy level of a BESS typically needs periodic balancing to preserve sufficient capacity for potential charging and discharging demands. Therefore, the inconvenience caused to a BESS user by operational adjustment can be reflected by the unbalanced energy at times of balancing.
@@ II-B3 Heat Pumps
We define the baseline power and accumulated energy consumption profiles, $\mathbf{p}^{\text{base}}$ and $\mathbf{e}^{\text{base}}$ , as those that always maintain the indoor temperature at the set point $\theta^{\text{set}}_{t},\forall t\in[T]$. Activating the flexibility of HPs leads to the fluctuation of indoor temperature within the upward and downward adjustment ranges relative to the temperature set point, denoted by $\Delta\hat{\theta}_{t}$ and $\Delta\check{\theta}_{t}$, respectively, (with vector forms $\Delta\hat{\boldsymbol{\theta}}\in\mathbb{R}^{T}_{+}$ and $\Delta\check{\boldsymbol{\theta}}\in\mathbb{R}^{T}_{+}$). Regarding the relationship between the temperature adjustment ranges, $\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\boldsymbol{\theta}}$, and the upward and downward adjustment ranges in power, $\Delta\hat{\mathbf{p}}$ and $\Delta\check{\mathbf{p}}$, and accumulated energy consumption, $\Delta\hat{\mathbf{e}}$ and $\Delta\check{\mathbf{e}}$, the following Proposition holds:
@@
$\exists\mathbf{D}\in\mathbb{R}^{T\times T}$, $\forall\Delta\hat{\boldsymbol{\theta}}\geq\mathbf{0}$ and $\Delta\check{\boldsymbol{\theta}}\geq\mathbf{0}$, if the indoor temperature profile $\boldsymbol{\theta}$ satisfies $-\Delta\check{\boldsymbol{\theta}}\leq\Delta\boldsymbol{\theta}-\boldsymbol{\theta}^{\text{set}}\leq\hat{\boldsymbol{\theta}}$, then the corresponding accumulated energy consumption profile $\mathbf{e}$ satisfies $-\Delta\check{\mathbf{e}}\leq\mathbf{e}-\mathbf{e}^{\text{base}}\leq\Delta\hat{\mathbf{e}}$, where
$\Delta\hat{\mathbf{e}}={\mathbf{D}}\Delta\hat{\boldsymbol{\theta}},\;\;\Delta\check{\mathbf{e}}={\mathbf{D}}\Delta\check{\boldsymbol{\theta}}.$ (6)
Specifically, $\mathbf{D}$ is an invertible matrix defined as:
${\mathbf{D}}\triangleq\frac{\Delta_{T}H}{\eta(1-\alpha)}\left[{\begin{array}[]{cccc}1&{}\hfil&{}\hfil&{}\hfil\\ {1-\alpha}&1&{}\hfil&{}\hfil\\ \vdots&\ddots&\ddots&{}\hfil\\ {1-\alpha}&\cdots&{1-\alpha}&1\end{array}}\right].$
However, $\nexists\mathbf{F}\in\mathbb{R}^{T\times T}$, such that $\forall\Delta\hat{\boldsymbol{\theta}}\geq\mathbf{0}$ and $\Delta\check{\boldsymbol{\theta}}\geq\mathbf{0}$, if $-\Delta\check{\boldsymbol{\theta}}\leq\boldsymbol{\theta}-\boldsymbol{\theta}^{\text{set}}\leq\Delta\hat{\boldsymbol{\theta}}$, then the corresponding power profile $\mathbf{p}$ satisfies $-\mathbf{F}\Delta\check{\boldsymbol{\theta}}\leq\mathbf{p}-\mathbf{p}^{\text{base}}\leq\mathbf{F}\Delta\hat{\boldsymbol{\theta}}$.
Proof: See Appendix.
Proposition 1 indicates that the upward and downward temperature adjustment ranges $\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\boldsymbol{\theta}}$ can be linearly transformed into the upward and downward adjustment ranges in accumulated energy consumption $\Delta\hat{\mathbf{e}}$ and $\Delta\check{\mathbf{e}}$, respectively. However, $\Delta\hat{\boldsymbol{\theta}}$ and $\Delta\check{\boldsymbol{\theta}}$ cannot be respectively transformed into upward and downward power adjustment ranges $\Delta\hat{\mathbf{p}}$ and $\Delta\check{\mathbf{p}}$. The rationale behind Proposition 1 can be intuitively understood: the indoor temperature is primarily influenced by the accumulated power over time rather than the power in each time slot itself.
The transformation in (6) allows aggregators to have a contract with HP users specifying the compensation for per unit temperature adjustment range and to transform this compensation into cost coefficients of energy adjustment range. We denote by $\hat{\boldsymbol{\rho}}\in\mathbb{R}^{T}_{+}$ and $\check{\boldsymbol{\rho}}\in\mathbb{R}^{T}_{+}$ the vectors of compensations for the activated upward and downward temperature adjustment ranges, respectively. Then, the total cost of the activated flexibility of an HP is:
$C_{\text{HP}}=\hat{\boldsymbol{\rho}}^{\top}\Delta\hat{\boldsymbol{\theta}}+\check{\boldsymbol{\rho}}^{\top}\Delta\check{\boldsymbol{\theta}},$ (7)
@@
$C_{\text{HP}}=\hat{\boldsymbol{\rho}}^{\top}\mathbf{D}^{-1}\Delta\hat{\mathbf{e}}+\check{\boldsymbol{\rho}}^{\top}\mathbf{D}^{-1}\Delta\check{\mathbf{e}},$ (8)
@@
Fig. 4 illustrates adjustment ranges in the indoor temperature, the corresponding adjustment ranges in accumulated energy consumption, and cost coefficients. Cost coefficients of energy adjustment ranges shown in this figure, which increase with time, are generated based on a constant cost coefficient of temperature adjustment range across the time horizon. Take $\check{\mathbf{c}}_{\text{e}}$ as an example, assuming that $\check{\rho}_{t}=\check{\rho}_{0},\forall t\in[T]$, then $\check{\mathbf{c}}_{\text{e}}=([\check{\rho}_{0},\check{\rho}_{0},\cdots,\check{\rho}_{0}]\mathbf{D}^{-1})^{\top}=\frac{\eta(1-\alpha)}{H\Delta_{T}}\check{\rho}_{0}[{\alpha^{T-1}},...,{\alpha^{2}},\alpha,1]^{\top}$. As $10<\alpha<$ 1 by definition, $\check{c}_{\text{e},t}$ increases with the time index $t$. Besides, $\check{\mathbf{c}}_{\text{e}}$ is typically larger than $\hat{\mathbf{c}}_{\text{e}}$ because downward temperature adjustments cause more discomfort to the people inside the building than upward adjustments considering the primary task of the HP is heating.
@@
Finally, we note that $\hat{\boldsymbol{\rho}}$ and $\check{\boldsymbol{\rho}}$ should not be uniform across all buildings, as the thermal dynamic parameters and HPs’ COPs vary. Identical temperature adjustments can lead to different energy adjustments in different buildings, so applying uniform $\hat{\boldsymbol{\rho}}$ and $\check{\boldsymbol{\rho}}$ would be unfair to HP users in a well-insulated building. Thus, $\hat{\boldsymbol{\rho}}$ and $\check{\boldsymbol{\rho}}$ should be customized based on the thermal dynamic parameters of the building and the HP’s COP. We suggest setting $\hat{\boldsymbol{\rho}}$ and $\check{\boldsymbol{\rho}}$ proportional to $\frac{H\Delta_{T}}{\eta(1-\alpha)}$, which aligns the HP’s flexibility cost coefficients with the contribution to energy adjustment ranges.
III Model and Costs of Aggregated Flexibility
We have discussed the flexibility cost formulation of individual DERs. This section focuses on the aggregator’s perspective, namely the model and cost for the aggregated flexibility, which the aggregators should report to the DSO. We denote the set of DERs in the aggregator as $\mathcal{K}$, indexed by $k$, and the power vector of the $k$-th DER as $\mathbf{p}_{k}\in\mathbb{R}^{T}$. Hence, the aggregated power is $\mathbf{P}\triangleq\sum\nolimits_{k\in\mathcal{K}}\mathbf{p}_{k}$. The aggregated flexibility model refers to the feasible region of $\mathbf{P}$. Given the individual flexibility constraints (2)-(4), the aggregated flexibility, as derived in our previous works [16] and [17], can be formulated as follows:
@@ III Model and Costs of Aggregated Flexibility
where $\mathbf{U}$ is an ${M\times T}$ coefficient matrix with each row vector being a ${1\times T}$ vector composed of zeros and ones, and $M$ represents the number of constraints that is dependent on whether an exact or an approximate model is used. Constraints (9) define a polytope in the $\mathbb{R}^{T}$ space with a specific shape. We let $\mathbb{B}\triangleq\{0,1\}$ and $\mathbb{B}^{T}$ denote the set of all $T$-dimensional row vectors composed solely of 0s and 1s. If the row vectors of $\mathbf{U}$ span $\mathbb{B}^{T}$ except for $\mathbf{0}$, i.e., $M=2^{T}-1$, $\mathbf{U}$ corresponds to the exact aggregation model. When the row vector set of $\mathbf{U}$ is a subset of $\mathbb{B}^{T}$, $\mathbf{U}$ corresponds to an approximate model. Bound parameters $\underline{\boldsymbol{\phi}}\in\mathbb{R}^{M}$ and $\overline{\boldsymbol{\phi}}\in\mathbb{R}^{M}$ are derived from all the individual power and energy bound parameters in (2)-(3). Reducing the number of rows of $\mathbf{U}$ creates an approximate model with low complexity, and appropriately adjusting $\underline{\boldsymbol{\phi}}$ and $\overline{\boldsymbol{\phi}}$ forms either an outer or inner approximation. For example, setting the coefficient matrix $\mathbf{U}$ to only include power and accumulated energy constraints and the bound parameters $\underline{\boldsymbol{\phi}}$ and $\overline{\boldsymbol{\phi}}$ to the sum of all individual power and energy bounds yields an outer approximation [16], and properly shrinking the gap between $\underline{\boldsymbol{\phi}}$ and $\overline{\boldsymbol{\phi}}$ forms an inner approximation [17]. Since the primary focus of this paper is on the quantification of flexibility costs and values rather than on flexibility aggregation modeling, we directly employ the inner approximated aggregation model proposed in our previous research [17], which has been validated for its advanced accuracy and computational efficiency.
We are now ready to introduce the flexibility cost formulation for the aggregated flexibility model (9). For the sake of clarity, we hereafter refer to the product of $\mathbf{U}$ and the aggregated power profile $\mathbf{P}$, i.e., $\mathbf{UP}$, as the flexibility trajectory and the product of $\mathbf{U}$ and the aggregated baseline power profile $\mathbf{P}^{\text{base}}$, i.e., $\mathbf{UP}^{\text{base}}$, as the baseline flexibility trajectory. Similar to the flexibility model (2)-(4) for individual DERs, activating the aggregated flexibility means that the flexibility trajectory $\mathbf{UP}$ can deviate within a certain range relative to the baseline flexibility trajectory $\mathbf{UP}^{\text{base}}$, which can be formulated as:
@@
where $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ in +$\mathbb{R}^{M}_{+}$ denote the upward and downward cost coefficients for flexibility adjustment ranges, respectively, which the aggregator reports to the DSO.
The next question is how the aggregator can calculate the cost coefficients $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$. Principally, $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ are related to the aggregator’s cost of purchasing flexibility from the DER users and its operation costs. Aggregators may develop various strategies to determine $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ for reporting to the DSO. Since the focus of this paper is primarily on the quantification of flexibility costs, we here use a simple approach to estimate $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ without a detailed discussion of the aggregators’ strategies.
First, it should be noted that the flexibility model (9) also includes constraints of power and accumulated energy consumption of the aggregator. In other words, the power and accumulated energy consumption are a part of the flexibility trajectory $\mathbf{UP}$. The power and accumulated energy part of cost coefficients $\hat{\mathbf{c}}_{\text{agg}}$ and $\check{\mathbf{c}}_{\text{agg}}$ are estimated via a weighted average method, with the weights being the ratio of the maximum power and energy adjustment ranges of each DER, i.e.,
@@
If the aggregated flexibility model (9) includes constraints beyond power and accumulated energy consumption, i.e., $M>|\mathcal{I}_{\text{p,e}}|$, the cost coefficients of adjustment ranges in those additional parts should be directly set to zero. This setting avoids the duplication of the cost calculation, as power and accumulated energy adjustments could lead to adjustments in other segments of the flexibility trajectory. This does not imply that other segments of the flexibility trajectory do not contribute to the flexibility cost of the aggregator since those segments are subject to constraints and indirectly influence the aggregator’s total flexibility cost $C_{\text{agg}}$. Indeed, for simplicity, the DSO can require aggregators to report a flexibility model with only power and accumulated energy constraints and to report the cost coefficients only for adjustment ranges in this segment. Nevertheless, since we have designed a more general flexibility cost formulation, it holds the potential to be applied to more accurate flexibility models.
IV The DSO Model for Flexibility Activation and Revenue Allocation
The DSO optimizes the activation of flexibilities from aggregators for energy arbitrage and ancillary service provision at the transmission level. Here, we primarily consider reserve capacities provision as the ancillary service for participation, while the proposed model also adapts to other ancillary services. In this context, the DSO should 1) determine the reference power profile and reserve capacities offered to the TSO, 2) determine the activated adjustment range for the flexibility trajectory of each aggregator, and 3) allocating the revenues obtained from the TSO to each aggregator.
IV-A The DSO Optimization Model for Flexibility Activation
The objective of the DSO is to minimize the net cost of the distribution system, formulated as: